### v1

# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

Abstract
In recent years, graph neural networks (GNNs) have become a popular tool to improve the accuracy and performance of recommender systems. Modern recommender systems are not only designed to serve the end users, but also to benefit other participants, such as items and items providers. These participants may have different or conflicting goals and interests, which raise the need for fairness and popularity bias considerations. GNN-based recommendation methods also face the challenges of unfairness and popularity bias and their normalization and aggregation processes suffer from these challenges. In this paper, we propose a fair GNN-based recommender system, called HetroFair, to improve items’ side fairness. HetroFair uses two separate components to generate fairness-aware embeddings: i) fairness-aware attention which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees, and ii) heterophily feature weighting to assign distinct weights to different features during the aggregation process. In order to evaluate the effectiveness of HetroFair, we conduct extensive experiments over six real-world datasets. Our experimental results reveal that HetroFair not only alleviates the unfairness and popularity bias on the items’ side, but also achieves superior accuracy on the users’ side. Our implementation is publicly available at https://github.com/NematGH/HetroFair
Keywords Recommendation systems, graph neural networks (GNNs), fairness, heterophily.
1 Introduction
Graph-based recommendation has become more popular in recent years, due to the existence of a bipartite graph in the nature of recommender systems [1, 2, 3, 4, 5, 6]. In this bipartite graph, nodes are divided into two sets users and items, and edges are ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [7, 8, 9]. With the widespread use of graph neural networks in recommender systems, the competition to build systems with higher accuracy entered a new phase [2, 10, 11, 12]. Early methods attempted to improve the accuracy of recommender systems, to better satisfy end users. However, users are not the only stakeholders and accuracy is not the sole objective. Recent researches demonstrate that recommender systems are multi-stakeholder systems in which in addition to end users, other participants (especially items and items providers) should also be considered, during the recommendation process [13, 14, 15, 16, 17]. On the other hand, the power law characteristic of the recommendation graphs shows that a large number of items possess a low degree, while only a few of them have a large degree. Figure 1 illustrates the degree distribution of two real-world recommendation datasets. This feature causes a phenomenon, known in the literature as popularity bias, where popular items are recommended more frequent than the other items. As a result, items with few interactions (long-tail items) are ignored, even if they are of high quality and relevance [18, 19]. Similar investigations have been done for GNN-based recommendation, revealing that long-tail items are often sacrificed during message normalization and aggregation of GNNs [20, 21]. Various methods have been proposed in recent years for long-tail recommendation, which aim to recommend items that belong to the long tail of the distribution of ratings or popularity and at the same time, try to achieve high accuracy on the users’ side. The authors of [22] introduce regularization techniques to overcome the popularity bias. Some other works concentrate on creating multiple graphs for different objectives and simultaneously optimize those graphs in an integrated manner [23, 24]. A few other works directly target the normalization and aggregation processes in the recommendation systems that utilize graph neural networks [20, 21]. However, these GNN-based recommenders either consider static weights for message normalization, or ignore different aspects of items/users. In this paper, we propose the HetroFair model, to improve items’ side fairness in GNN-based recommendation systems. To do so, first we directly address the cause of popularity bias and reduce the effect of nodes’ degrees on enlarging the values of their embeddings, via involving dot products during the normalization process. Second, with respect to the heterophily property in graphs, we distinguish between different aspects of an item, by considering an specific weight for each feature during the aggregation process. In the heterophily property, nodes of the graph that are not similar to each other, tend to be connected to each other. To the best of our knowledge, our model is the first approach that addresses fairness in GNN-based recommendation using the heterophily property. Our key contributions are summarized as follows:
  • We theoretically analyze the effect of message normalization in GNN-based recommendation on popularity bias.
  • To generate fair embeddings for users and items, we design a fairness-aware attention mechanism and extend this attention to each feature by considering the heterophily property in graphs.
  • We propose a novel GNN-based recommendation system, called HetroFair, that exploits heterophily to improve fairness.
  • We conduct extensive experiments over six well-known datasets and show that our proposed model outperforms state of the art methods, in terms of both accuracy and fairness metrics. Furthermore, by performing ablation studies, we demonstrate that each component of our proposed method considerably contributes in the performance of the method.
The rest of this paper is organized as follows. In Section 2, we provide a brief overview of related methods. In Section 3, we introduce necessary preliminaries and definitions. In Section 4, we describe our proposed method in details. In Section 5, we report the results of our extensive experiments conducted to show the high performance of our proposed method. Finally, the paper is concluded in Section 6.
(a) Epinions
(b) CDs
Figure 1: Degree distribution of items, in two recommendation datasets.
2 Related work
In this section, we review recent advances in two fields that are closely related to our work: graph-based recommendation and fairness-aware recommendation.
2.1 Graph-based recommendation
After the distinct capability of GNNs in node classification within graph-structured data [25, 26, 27], the GC-MC method [1] extends this model to link prediction. It utilizes a graph auto-encoder to fill the user-item interaction matrix. PinSage [28] introduces a recommender system for large-scale graphs, leveraging the inductivity property of graph neural networks, a concept proposed by GraphSage [26]. NGCF [4] models higher-order connectivities by combining embeddings from different layers after embedding propagation on the user-item bipartite graph. LR-GCCF [29] removes non-linearity and demonstrates that this removal enhances recommendation performance. Taking it a step further, LightGCN [2] removes non-linearity and weight transformation, introducing a lightweight version of graph convolution for embedding propagation in user-item bipartite graphs. UltraGCN [11], bypasses explicit message passing. Instead, it aims to compute the limit of the infinite graph convolution layer by leveraging the constraint loss. IGCCF [10] introduces an inductive convolution-based algorithm for recommender systems without training individual user embeddings, by leveraging an item-item graph. Despite the high performance of these graph-based recommendation systems in improving accuracy, they often neglect to investigate unfairness resulting from popularity bias, and fail to use techniques that simultaneously consider items’ side.
2.2 Fairness-aware recommendation
Research in the domain of fairness in graph-based recommendation is divided into two categories: diverse recommendation and popularity debiasing.
Diverse recommendation
Diverse recommendation means that the items in individual recommendation lists should be dissimilar and cover a broader range of categories. BGCF [30] utilizes the node-copy mechanism and connects users to higher-order item neighbors that belong to different categories than their direct neighbors, yet remain similar. DGCN [31] mitigates the influence of dominant categories while elevating the significance of disadvantaged categories within neighboring nodes, achieved through re-balanced neighborhood sampling. Additionally, it leverages adversarial learning to achieve category-free representation for items. Isufi et al. [32] suggest jointly learning from nearest and furthest neighbor graphs to achieve accuracy-diversity trade-off for node representations. DGRec [33] applies a sub-modular function for the neighbor selection process to recommend a diverse subset of items. It also manipulates the loss function, increasing the loss value for long-tail items and, conversely, decreasing it for short-head items.
Popularity debias
The primary objective of popularity debias is to unearth qualified but less interacted items, long-tail ones, which are items that have lower training samples than short-head items. V2HT [23] constructs a heterogeneous item-item graph with four different types of edges to connect long-tails to popular items. SGL [34] brings self-supervised learning to the recommendation setting and creates an auxiliary task for the long-tail recommendation. r-AdjNorm [20] changes the power of the symmetric square root normalization term in graph neural networks to control the normalization process during neighborhood aggregation, aiming for improved results, especially for low-degree items. APDA [21] assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. Our proposed method takes a different approach to improve items’ side fairness and on the one hand, considers different weights for features during the aggregation process. On the other hand, it directly involves dot products in the normalization process.
3 Preliminaries
In this section, we introduce preliminaries widely used in the paper, including task formulation, graph neural networks, the Light Graph Convolution model and the heterophily property of graphs.
3.1 Problem setup
In graph-based recommendation, we have a set of users $U=\{u_{1},u_{2},\ldots,u_{|U|}\}$ and a set of items $I=\{i_{1},i_{2},\ldots,i_{|I|}\}$ that constitute the nodes $V$ of a bipartite graph $G=\left(V,E\right)$, where $V=U\cup I$ and $E$ represents implicit interactions (such as purchases, views and clicks) between users and items. The interactions matrix $R\in\mathbb{R}^{|U|\times|I|}$ is defined such that if $e_{u,i}\in E$ then $R_{u,i}=1;$ otherwise, it is 0. In graph $G$, there exist no edges connecting users or items to each other. Thus, the graph adjacency matrix $A$ is denoted as:
$A=\begin{bmatrix}0_{|U|\times|U|}&\hskip 1.0pt\vrule\hskip 1.0pt&R\\ \hline\cr R^{T}&\hskip 1.0pt\vrule\hskip 1.0pt&0_{|I|\times|I|}\\ \end{bmatrix}.$
This matrix is used during message propagation by a graph neural network.
3.2 Graph neural networks
Graph neural networks (GNNs) are a class of learning algorithms that work well on graph data structures. The primary task of a GNN is to map each node in the graph to a vector in a low-dimensional vector space. This vector is called the embedding or representation of the node. GNNs rely on message passing, wherein each node in the graph sends a message to its direct neighbors. Messages, in this context, are intermediate vector representations of nodes. Graph convolutional network (GCN) [25] is one of the first and most popular graph neural networks that uses the following message passing rule:
$H^{\left(k+1\right)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{\left(k\right)}M^{\left(k\right)}\right).$ (1)
Here, $D$ is the diagonal degree matrix that has the same dimension as the adjacency matrix $A$ of the graph, $D_{ii}=\sum_{j}A_{ij}$, $\tilde{D}=D+I$, $\tilde{A}=A+I$, $I\in R^{\left(|U|+|I|\right)\times\left(|U|+|I|\right)}$ is the identity matrix, $\sigma\left(\cdot\right)$ is an activation function, $H^{\left(k\right)}$ is the matrix of node embeddings (representations) in the layer $k$ and $M^{\left(k\right)}$ denotes a layer-specific trainable weight matrix.
3.3 Light Graph Convolutional
The authors of LightGCN [2] demonstrate that non-linearity and feature transformation in the standard GCN are unnecessary for collaborative filtering graphs. They introduce a light version of GCN, as follows:
$h_{u}^{\left(k+1\right)}=\sum_{i\in N(u)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}h_{i}^{\left(k\right)},$ (2)
$h_{i}^{\left(k+1\right)}=\sum_{i\in N(i)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}h_{u}^{\left(k\right)}\cdot$ (3)
Here, $h_{u}^{\left(k\right)}$ and $h_{i}^{\left(k\right)}$ denote the embeddings of user $u$ and item $i$ at layer $k$. By $N\left(x\right)$ we denote the set of nodes which have an edge to node $x$, and by $d_{x}$ we denote the degree of node $x$. The term $\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}$ is called the symmetric square root normalization term. In the matrix form, $H^{\left(0\right)}=[h_{1}^{\left(0\right)},h_{2}^{\left(0\right)},\ldots,h_{|V|+|E|}^{\left(0\right)}]$ denotes the initial embeddings layer. The embeddings in the next layers are formulated as follows:
$H^{\left(k+1\right)}=(D^{-\frac{1}{2}}AD^{-\frac{1}{2}})H^{\left(k\right)}.$ (4)
3.4 Heterophily in graphs
Heterophily refers to a property in which nodes of a graph that are not similar to each other, tend to be connected to each other. In other words, it describes the preference for forming connections with nodes that have different characteristics or attributes. This concept is contrasted with homophily, where nodes with similar attributes are more likely to be connected. The homophily rate is a measure used to quantify the tendency of nodes to connect with other nodes that have similar attributes. It can be computed using the following formula [35]:
homophily $\text{homophily rate}=\frac{1}{|V|}\sum_{v\in V}\frac{|\{(w,v):w\in N(v)\wedge C_{v}=C_{w}\}|}{d_{v}},$ (5)
where $C_{x}$ is the class label of node $x$. The range of this metric is within $[0,1]$, where a value of 1 indicates strong homophily, meaning that neighbors between two directions of edges have similar features, in terms of node labels. A value of 00, on the other hand, signifies that neighbor nodes have diverse characteristics [35].
4 Our proposed method
In this section, we first analyze the effect of symmetric square root normalization on popularity bias in GNN-based recommendation systems. We then present our approach to overcome this bias. Subsequently, we outline the process of training our recommendation model. Lastly, we provide the pseudocode of our proposed method and analyze its time complexity.
4.1 Symmetric square root normalization effect
The following theorem analyzes the over-smoothing problem in graph neural networks. Our research investigates the implications of this foundational result in the context of fairness in graph-based recommendation systems.
Theorem 1
[36] For every undirected connected graph $G=\left(V,E\right)$ where each node has a self loop, the following holds:
$\small\lim_{{k\to\infty}}\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}\right)_{i,j}^{k}=\frac{\sqrt{\left(d_{i}+1\right)\left(d_{j}+1\right)}}{2|E|+|V|}.$ (6)
With respect to Light Graph Convolution in Equation 4, after $k$ times message propagation, the last layer node embeddings can be written as:
$H^{\left(k\right)}=\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}\right)^{k}H^{\left(0\right)}.$ (7)
For user $u$, Equation 7 can be written as follows:
$h_{u}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}}{2|E|+|V|}\left[\left(\sqrt{d_{1}+1}\right)h_{1}^{\left(0\right)}+\left(\sqrt{d_{2}+1}\right)\right.\\ \left.h_{2}^{\left(0\right)}+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right].$ (8)
In recommendation systems, dot product is a commonly used operator to identify items that are relevant to users. Let $i$ and $i^{\prime}$ be two items that exist in the test profile of user $u$. The prediction scores $y_{ui}$ and $y_{ui^{\prime}}$ are calculated as follows:
$y_{ui}=h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2},$ (9)
$y_{ui}=h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)}=\frac{\sqrt{d_{u}+1}{\sqrt{d_{i^{\prime}}+1}}}{\left(2|E|+|V|\right)^{2}}\left[\sqrt{\left(d_{1}+1\right)}h_{1}^{\left(0\right)}\right.\\ \left.+\ldots+\left(\sqrt{d_{|V|}+1}\right)h_{|V|}^{\left(0\right)}\right]^{2}.$ (10)
The only difference between Equations 9 and 10 is the subscripts in the terms $\sqrt{d_{i}+1}$ and $\sqrt{d_{i^{\prime}}+1}$ in the numerator of the Equations. Therefore, if the degree of node $i$ surpasses that of node $i^{\prime}$, it leads to a higher score for node $i$, potentially neglecting the users’ preference for low-degree nodes, despite our knowledge that the user likes both of them. This observation highlights a flaw in the prevalent symmetric square root normalization, employed by graph neural networks, as it tends to exhibit popularity bias and sacrifices items with lower degrees. In an effort to rectify this bias and alleviate its impact on ranking, we propose a novel fairness-aware attention-based model. This model aims to enhance the ranking opportunities for long-tail items, ensuring they have a fair chance of attaining higher positions.
4.2 Fairness-aware attention
In the previous section, we demonstrated the destructive effect of symmetric square root normalization on the representation of long-tail items. To address this issue, we change light graph convolution, proposed by LightGCN [2], as follows:
$h_{u}^{\left(k\right)}=\sum_{i\in N\left(u\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\lambda\times\sigma(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)})}h_{i}^{\left(k-1\right)},$ (11)
$h_{i}^{\left(k\right)}=\sum_{u\in N\left(i\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\lambda\times\sigma(h_{i}^{\left(k-1\right)}\cdot h_{u}^{\left(k-1\right)})}h_{u}^{\left(k-1\right)},$ (12)
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. By incorporating the sigmoid of dot products alongside the previous normalization term, messages from long-tail items (such as $i$) receive a higher importance coefficient than short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients. This not only prevents the over-enlargement of representations but also ensures that the effects of large representations are gradually transmitted through the computational graph neighborhood.
4.3 Heterophily feature weighting
In recommender systems, users and items exhibit diverse characteristics. For instance, in movie recommender systems, individual movies, as items, possess unique attributes. Movies vary in terms of cast members, with differing numbers of renowned actors featured. Additionally, movies belong to distinct genres, and some may even encompass multiple genres. Consequently, we exploit homophily rate to understand users’ preferences for these diverse items’ characteristics. To achieve this goal, we construct an item-item graph, linking items within a users’ profile. Each item is assigned its category as a label. Table 1 summarizes the item-item graph data statistics [^1] In Section 5, we use six datasets for our empirical evaluations. However, in Table 1, only five datasets are reported. The reason is that the Epinions dataset lacks the category information necessary for constructing an item-item graph. Therefore, we do not report it in Table 1.. The homophily rate is determined using the GEOM-GCN method [35].
Table 1: Item-item graph statistics.
Statistics #Nodes #Edges #Classes Homophily
Amazon-Electronics 11589 3790708 528 0.03
Amazon-Beauty 733 119329 112 0.05
Amazon-Health 1260 379736 290 0.01
Amazon-Movie 21901 43758979 77 0.7
Amazon-CDs 16184 18113968 209 0.11
We notice from Table 1 that:
  • Low homophily rate indicates that the nodes in the graph have diverse attributes and characteristics. In essence, there is a limited tendency for nodes to connect with others sharing similar features. This observation indicates a high level of diversity within the network.
  • Nodes may form connections based on a variety of factors, and the network might be characterized by a mixture of different classes or categories. This diversity could be beneficial for information diffusion or collaboration across various domains.
Based on the above analysis, instead of assigning a unique weight $\sigma(h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)})$ to the message between $u$ and $i$, we opt for distinct weights for different features in each propagated message. This enables us to effectively capture dissimilarities between user and item features. So, we define the hererophily feature weighting as follows:
$w_{ui}^{\left(k\right)}=\lambda\times\sigma\left(\left(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)}\right)W^{\left(k\right)}\right),$ (13)
where $W^{\left(k\right)}\in\mathbb{R}^{1\times d}$ is a trainable linear transformation, $d$ is the dimension of node embeddings and $w_{ui}^{(k)}\in\mathbb{R}^{1\times d}$. Therefore, Equations 11 and 12 are revised as follows:
$h_{u}^{\left(k\right)}=\sum_{i\in N\left(u\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}h_{i}^{\left(k-1\right)}\oslash w_{ui}^{\left(k\right)},$ (14)
$h_{i}^{\left(k\right)}=\sum_{u\in N\left(i\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}}h_{u}^{\left(k-1\right)}\oslash w_{ui}^{\left(k\right)},$ (15)
[…]
4.4 Model training
To optimize model parameters, we use the well known pairwise Bayesian Personalized Ranking (BPR) loss function [37]. This loss function is designed to predict higher rank scores for items that users interact with, compared to unobserved items:
$L_{BPR}=\sum_{\left(u,i,j\right)\in T}-\ln\sigma\left(\tilde{y}_{ui}-\tilde{y}_{uj}\right)+\beta||\Theta||^{2},$ (16)
where $O_{u}^{+}$ is user intracted items during training and $T=\left\{\left(u,i,j\right)\mid i\in O_{u}^{+}\wedge j\in I\setminus O_{u}^{+}\right\}$ indicates the training set. The hyperparameter $\beta$ controls regularization to prevent overfitting, $||\Theta||$ is the model’s parameters that must be learned during the optimization process, $\tilde{y}_{ui}$ and $\tilde{y}_{uj}$ are user interest scores for items $i$ and $j$ respectively, obtained through the dot product of the embeddings of the user and item.
4.5 The algorithm
In this section, we present the process of fair embedding generation, along with the algorithmic steps that our recommendation algorithm follows during a batch. Algorithm 1 describes one batch training process, in our model. First, in Lines 3-6, we initialize model parameters using the Xavier initialization method [38]. The embedding matrix, updated during the training process, is also initialized. In Line 9, we call the FairEmbeddingGeneration function, presented in Algorithm 2 to calculate users’ and items’ embeddings. Subsequently, for every user-item pair in the current batch $O_{batch}$, we choose a negative sample uniformly at random with replacement, as explained in [37]. In Lines 12-15, we compute the interest score of the user for positive and negative items, calculate the loss, and finally backpropagate the error to update model parameters. Algorithm 2 outlines the fair embedding generation process. For every node in the graph and for each layer, we act as the follow. First in Line 7, we compute the fairness-aware attention as the dot products of the embeddings of users and items, for all items in the users’ training profile. Then, we transform these attention weights into a new vector space to consider distinct attention for each feature, as shown in Line 8. In Line 10, by taking the inverse of the weight in the message passing process, we enhance the likelihood of suggesting long-tail items to users while simultaneously diminishing the dominance of popular items. Finally, in Line 12, we combine the embeddings from different layers to obtain the final representations of items/users.
Algorithm 1 One batch training process.
1:Input: Graph $G\left(V,E\right)$, the number of layers $K$.
2:Output: Updated model parameters.
3:$X\leftarrow$ Xavier initialization $\triangleright$ Initialize input features
4:for $k=1$ to $K$ do
5:     $W^{\left(k\right)}\leftarrow$ Xavier Initialization $\triangleright$ Initialize weight matrices
6:end for
7:$H\leftarrow X$triangleright Initialize embedding matrix
8:$L\leftarrow 0$triangleright Initialize BPR loss
9:$H\leftarrow$ FairEmbeddingGeneration($G,H,K$)
10:for $\left(u,i\right)\in O_{batch}$ do $\triangleright$ $O_{batch}$ is the set of $\left(u,i\right)$ pairs in the current batch
11:     Sample a negative item $j$ uniformly at random from $I\setminus O_{u}^{+}$
12:     $\tilde{y}_{ui}\leftarrow H\left[u\right]\cdot H\left[i\right]$
13:     $\tilde{y}_{uj}\leftarrow H\left[u\right]\cdot H\left[j\right]$
14:     $L\leftarrow L+\left(-\ln\sigma\left(\tilde{y}_{ui}-\tilde{y}_{uj}\right)+\beta||\Theta||^{2}\right)$
15:     Backpropagate and update model parameters for every node in $G$, using gradient descent on $L$
16:end for
17:return $\Theta$
Algorithm 2 FairEmbeddingGeneration
1:Input: Graph $G\left(V,E\right)$; input features matrix $X\in\mathbb{R}^{|V|\times d}$; number of layers $K$; weight matrices $W^{\left(k\right)}\in\mathbb{R}^{1\times d},\forall k\in\{1,\ldots,K\}$; parameter $\lambda$$\cdot$
2:Output: Matrix representation $Z\in\mathbb{R}^{|V|\times d}$ where each row in the matrix is the representation of node $v,\forall v\in V$.
3:$h_{v}^{(0)}\leftarrow x_{v},\forall v\in V$
4:for $v\in V$ do
5:     for $k=1$ to $K$ do
6:         for $i\in N(v)$ do
7:              $s\leftarrow h_{v}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)}$
8:              $w_{vi}^{\left(k\right)}\leftarrow\lambda\times\sigma(\begin{bmatrix}s\end{bmatrix}\times W^{\left(k\right)})$triangleright By $\begin{bmatrix}s\end{bmatrix}$, we mean the $1\times 1$ matrix, consisting of scalar $s$.
9:         end for
10:         $h_{v}^{\left(k\right)}\leftarrow\frac{1}{\sqrt{d_{v}}}\cdot\sum_{i\in N(v)}\frac{1}{\sqrt{d_{i}}}\cdot h_{i}^{\left(k-1\right)}\oslash w_{vi}^{(k)}$
11:     end for
12:     $Z\left[v\right]\leftarrow\frac{1}{K}\sum_{k=0}^{K}h_{v}^{\left(k\right)}$
13:end for
14:return $Z$
Time complexity
It is imperative for recommendation systems to deliver not only satisfactory performance but also operate efficiency. Striking a balance between these two criteria is often viewed as a trade-off. In this context, we delve into the analysis of the time complexity of our proposed model. If we neglect Lines 6-9 of Algorithm 2, our model becomes the same as the LightGCN model, which has the time complexity of $O\left(|E|Kd\right)$ in the graph convolution part, where $K$ and $d$ are constants [39]. Thanks to parameter sharing, the primary operation in Equation 13, which appears in Lines 7-8, is the multiplication of a $1\times 1$ matrix with a $1\times d$ matrix, which can be done in $O\left(d\right)$ time. Hence, the additional time complexity that our proposed method imposes on the basic LightGCN model is $O\left(d\right)$. According to this analysis, the time complexity of our proposed graph convolution process is $O\left(|E|\left(Kd+d\right)\right)=O\left(|E|Kd\right)$, which is the same as the LightGCN model.
5 Experiments
In this section, we report the results of our extensive experiments, conducted over six real-world datasets, and demonstrate the high accuracy and performance of our proposed method.
5.1 Experimental setup
5.1.1 Datasets
We examine the algorithms over six well-known datasets: Epinions [^2] https://snap.stanford.edu/data/soc-Epinions1.html [21], Amazon-Movies [^3] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html [40], Amazon-Electronics3 [40], Amazon-Health3 [40], Amazon-Beauty3 [40] and Amazon-CDs3 [40]. For the amazon review datasets, we adopt a $10$-core setting, filtering out users and items with less than $10$ interactions. In the case of Epinions, we utilize the data released by APDA  [21]. Table 2 depicts the statistics of the datasets.
5.1.2 Baselines
In order to evaluate the performance of HetroFair, we consider the following state-of-the-art recommendation models:
  • LightGCN [2]: This model simplifies graph convolution by excluding non-linearity and weight transformation. The final embedding is computed through a weighted sum of embeddings learned across various layers. Following the original paper, we use $a_{k}=\frac{1}{K+1}$ as the coefficient of layer $k$, during layer combination.
  • r-AdjNorm [20]: This model adjusts the strength of the normalization term to control the normalization process during neighborhood aggregation, aiming to improve results especially for low-degree items. Following the advice in the paper, we fine-tune the parameter $r$ within the range $[0.5,1.5]$ with a step size of $0.05$.
  • APDA [21]: This method assigns lower weights to connected edges during the aggregation process and employs residual connections to achieve unbiased and fair representations for users and items in graph collaborative filtering. Following the approach outlined in the original paper, we fine-tune the residual parameter $\lambda$ within the range $\left[0,1.0\right]$.
Table 2: Datasets statistics.
Datasets #User #Item #Intraction Density
Electronics 20242 11589 347393 0.148%
Epinions 11496 11656 327942 0.245%
Beauty 1340 733 28798 2.93%
Health 2184 1260 55076 2%
Movies 33326 21901 958986 0.131%
CDs 15592 16184 445412 0.176%
5.1.3 Evaluation metrics
Fainess metrics
There are two distinct approaches to examine fairness on the items’ side in recommender systems:
  • Statistical Parity (or Demographic Parity): this approach necessitates that different groups of items or similar individuals are treated similarly, without considering users’ interests. This approach fails to consider the users’ side and increases the risk of damaging users’ satisfaction.
  • Equal Opportunity: this approach mandates that groups of items benefiting from the outcome are afforded equal opportunities. It strives to treat all items fairly within users’ profiles, regardless of their popularity.
Researchers in [22] introduce two metrics based on equal opportunity:
$PRU=-\frac{1}{|U|}\sum_{u\in U}SRC\left(d_{\tilde{O}_{u}^{+}},rank_{u}\left(\tilde{O}_{u}^{+}\right)\right),$ (17)
$PRI=-\sum_{i\in I}SRC\left(d_{i},\frac{1}{U_{i}}\sum_{u\in U_{i}}rank_{u}\left(i\right)\right)\cdot$ (18)
Here, $SRC\left(\cdot,\cdot\right)$ calculates Spearman’s rank correlation, $\tilde{O}_{u}^{+}$ represents items in the test profile of user $u$, $rank_{u}\left(x\right)$ signifies the ranking of item $x$ that the model predicts for user $u$, $U_{i}$ is a set of users that item $i$ already exists in their test profiles and $d_{x}$ is the degree of node $x$. While PRU investigates items’ position bias for individual users’ recommendation lists, PRI considers the items’ side fairness across all users. A large value of the latter may result in long-tail items having limited opportunities to become popular [22]. In general, the methods’ performance in promoting fairness improves as the values of these two criteria get closer to 00.
Accuracy metrics
As ranking holds paramount importance in recommender systems, we go beyond conventional machine learning metrics and incorporate rank-aware metrics to evaluate the performance of our model. Specifically, we employ MRR (mean reciprocal rank), MAP (mean average precision), and NDCG (normalized discounted cumulative gain) as accuracy metrics. In the following, we provide a brief description of each metric. NDCG denotes the normalized discounted cumulative gain and $NDCG@N$ for top $N$ recommended items is defined as follow:
$\small NDCG@N=\frac{DCG@N}{IDCG@N},$ (19)
where $DCG@N$ and $IDCG@N$ are defined as follows:
$DCG@N=\sum_{i=1}^{N}\frac{2^{r\left(i\right)}-1}{log_{2}(i+1)},$ (20)
$IDCG@N=\sum_{i=1}^{REL_{N}}\frac{2^{r\left(i\right)}-1}{log_{2}(i+1)}.$ (21)
Here, $REL_{N}$ is a sorted list of the $N$ most relevant items in ascending order, $r(i)$ is the relevance of the $i^{th}$ ranked item, which is either 00 or $1$ respectively for irrelevant and relevant items. MRR is concerned with the position of the first relevant item:
$MRR=\frac{1}{|U|}\sum_{u\in U}\frac{1}{first_{u}},$ (22)
where $first_{u}$ refers to the position of the first relevant item. MAP is computed as the mean of average precision:
$MAP=\frac{1}{|U|}\sum_{u\in U}AP_{u},$ (23)
where $AP_{u}$ is the area under precision-recall curve. $AP$ is calculated using the following formula:
$AP=\frac{1}{number\ of\ relevant\ documents}\sum_{i=1}^{N}P@i\cdot r(i),$ (24)
where $P@i$ is precision at $i$.
5.1.4 Hyper-parameters setting
For all the methods, we use the same early stopping mechanism, stopping training after $15$ epochs if NGCF does not show improvement in these $15$ epochs. The embeddings dimension is set to $64$ for LightGCN and AdjNorm, $256$ for the APDA model, and $128$ for our model. The only exception is the Movies dataset where due to hardware limitations, we set the embeddings dimension to 64. The batch size for LightGCN and AdjNorm is set to $1024$, as proposed in their papers, while for APDA we set it to $2048$. The learning rate for all the baselines is set to $0.001$. For all the methods, we set both the regularization hyper-parameter $\beta$ and the weight decay parameter to $0.0001$. The learning rate for our model is set to $0.0005$. We set the number of propagation layers to $3$ in LightGCN and r-AdjNorm, and to $4$ in APDA and HetroFair.
5.2 Performance comparison
Table 3 shows the results of HetroFair in comparison to the other methods. The findings indicate that our method outperforms the other methods, not only in the fairness metrics but also in the accuracy metrics. These results show that our method is better fitted with the data and takes into account their distinct features.
Table 3: Performance comparison of the models, over the used datasets.
Dataset Metric LightGCN r-AdjNorm APDA HetroFair w/o Hetro
Epinions NDCG 0.0789 0.0787 0.0859 0.0895 0.0889
MRR 0.1361 0.1347 0.1471 0.1525 0.1522
MAP 0.0329 0.0328 0.0366 0.0379 0.0377
PRU 0.5233 0.5310 0.5188 0.4706 0.5054
PRI 0.5129 0.5404 0.4983 0.4384 0.4573
Electronics NDCG 0.0463 0.0440 0.0517 0.0525 0.0488
MRR 0.0645 0.0608 0.0708 0.0733 0.0676
MAP 0.0221 0.0207 0.0250 0.0256 0.0232
PRU 0.5216 0.5469 0.5548 0.4600 0.5020
PRI 0.5417 0.5834 0.5692 0.4462 0.4805
CDs NDCG 0.1231 0.1219 0.1379 0.1449 0.1346
MRR 0.1715 0.1609 0.1940 0.2017 0.1882
MAP 0.0611 0.0588 0.0710 0.0747 0.0680
PRU 0.3304 0.3366 0.3265 0.2972 0.3266
PRI 0.2824 0.3297 0.2763 0.2264 0.2495
Health NDCG 0.1259 0.1283 0.1314 0.1334 0.1351
MRR 0.1986 0.2040 0.2073 0.2112 0.2093
MAP 0.0595 0.0614 0.0637 0.0656 0.0669
PRU 0.5162 0.5330 0.5205 0.4922 0.4830
PRI 0.2164 0.2088 0.2266 0.1821 0.1270
Beauty NDCG 0.2113 0.2184 0.2274 0.2308 0.2302
MRR 0.2525 0.2616 0.2790 0.2824 0.2798
MAP 0.1200 0.1267 0.1354 0.1364 0.1389
PRU 0.4428 0.4221 0.4418 0.4139 0.3117
PRI 0.2164 0.2088 0.2266 0.1821 0.1270
Movies NDCG 0.0672 0.0681 0.0715 0.0777 0.0741
MRR 0.0966 0.0982 0.1011 0.1093 0.1048
MAP 0.0304 0.0308 0.0331 0.0365 0.0340
PRU 0.3098 0.3404 0.3384 0.3045 0.3251
PRI 0.3951 0.4545 0.3862 0.3355 0.3590
5.3 Ablation study
In this section, we explore the impact of different components of our proposed model on its performance. More precisely, we compare HetroFair against the situation wherein the heterophily feature weighting component is not used. We refer to this case as w/o Hetro. The results of our analysis are available in Table 3.
  • When we discard the heterophily feature weighting component, over the datasets Epinions, Electronics, CDs and Movies, all the metrics deteriorate. This observation indicates that the heterophily feature weighting component effectively improves both fairness and accuracy metrics.
  • Over the Health and Beauty datasets, w/o Hetro achieves better results than HetroFair. One possible explanation for this phenomenon is over-smoothing. In the case of w/o Hetro, our experiments are conducted using 2 GNN layers, which is the optimal value for it. However, we run HetroFair with 4 GNN layers, which is its default value. The diameter of the Health and Beauty datasets is 8, a very small value to set the number of GNN layers to 4. It is known that in order to avoid over-smoothing in graph neural networks, the number of GNN layers should be considerably less than the diameter of the graph dataset. When for these two datasets we decrease the number of GNN layers from 4 to 2, the performance of HetroFair significantly improves and it outperforms w/o Hetro. Over the Health dataset, the PRU and PRI metrics drop from $0.4922$ and $0.5028$ to $0.4622$ and $0.4794$, respectively, outperforming PRU and PRI of w/o Hetro. The accuracy metrics also increase, for example, NDCG increases from $0.1334$ to $0.1380$. The same trends hold for the Beauty dataset.
5.4 Hyper-parameters study
In this section, we investigate the effect of different hyper-parameters on models’ performance. Specifically, we study the impact of three hyper-parameters: the number of layers $K$, the embeddings size $d$ and $\lambda$. Due to the variety of the used datasets, we study each of the parameters over two datasets.
Number of layers
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN model drastically declines after 4 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
Figure 2: Comparing the models over the Epinions dataset, for different values of embeddings dimension.
Figure 3: Comparing the models over the CDs dataset, for different values of embeddings dimension.
Figure 4: The performance of different models, over the Epinions dataset, for various numbers of message passing layers.
Figure 5: The performance of different models, over the Beauty dataset, for various numbers of message passing layers.
Embeddings dimension
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 2 and 3. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. In Figures 1(a) and 2(a), the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity. In the investigation of the baseline models, it can be pointed out that no model consistently outperforms the other baselines and by revising the parameters, the relative performances of the models change with respect to each other. For example in Figure 3, we see that in most cases, LightGCN achieves better fairness scores than APDA, suggesting that the simple LightGCN method effectively captures data characteristics, particularly when a larger embeddings dimension is used. On the other hand, r-AdjNorm has the worst performance in fairness metrics, indicating that it fails to make a trade-off between accuracy and fairness.
$Parameter\lambda$
Parameter $\lambda$ plays a crucial role in the accuracy-fairness trade-off. We consider the datasets Health and Beauty, and change the value of $\lambda$ from $0.05$ to $0.95$ with $0.05$ step. The results are depicted in Figures 6 and 7. Over both datasets, we observe almost the same pattern: within first few steps, the fairness metrics are in their lowest values, while the accuracy metrics are also too low. As the value of $\lambda$ increases, the performance of HetroFair increases, while PRU and RRI deteriorate until reaching to an optimal value for $\lambda$, which leads to almost stable values for all the metrics.
Figure 6: Studying the effect of $\lambda$, over the Beauty dataset.
Figure 7: Studying the effect of $\lambda$, over the Health dataset.
5.5 Short-head and long-tail items performance
To evaluate the effectiveness of HetroFair in recommending long-tail items, we modify the prior test set and construct two following test sets, one for long-tail items and the other for short-head items. For the long-tail test set, we exclude all short-head items from the $\tilde{O}_{u}^{+}$ sets of all users $u$. For the short-head test set, we remove all long-tail items existing in the $\tilde{O}_{u}^{+}$ sets. We sort the items based on their degree in descending order and consider the top $20\%$ of them as short-head items, treating the rest as long-tail items. Figure 8 depicts the results of this investigation, over all the six datasets. These figures illustrate that for the category of long-tail items, in most cases HetroFair achieves better accuracies. This is compatible with the results presented in Table 3, which indicate a reduction in the fairness metrics. Moreover, this improvement coincides with an increase in accuracy in the category of short-head items. Combining the results of Table 3 with the results presented in Figure 8 depicts that our performance improvement is not solely tied to long-tail items. The simultaneous improvement of HetroFair in recommending short-head items also plays a substantial role in reinforcing its overall performance.
(a) Long-tail items performance on the Epinions dataset.
(b) Short-head items performance on the Epinions dataset.
(c) Long-tail items performance on the Electronics dataset.
(d) Short-head items performance on the Electronics dataset.
(e) Long-tail items performance on the CDs dataset.
(f) Short-head items performance on the CDs dataset.
(g) Long-tail items performance on the Health dataset.
(h) Short-head items performance on the Health dataset.
(i) Long-tail items performance on the Beauty dataset.
(j) Short-head items performance on the Beauty dataset.
(k) Long-tail items performance on the Movies dataset.
(l) Short-head items performance on the Movies dataset.
Figure 8: Performance comparison of the methods, for short-head and long-tail items.
6 Conclusion
In this paper, we studied items’ side fairness in GNN-based recommendation systems. First, we pointed out that message normalization and aggregation processes in GNNs could lead to popularity bias. Then, we proposed the HetroFair model, that utilizes two techniques to overcome this bias and generate fair embeddings. The first technique was fairness-aware attention, which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees. The second technique was heterophily feature weighting, which assigns distinct weights to different features during the aggregation process. Our extensive experiments over six well-known datasets demonstrated the high performance of HetroFair, in improving both users’ side accuracy and items’ side fairness.